# Integral canonical models of exceptional Shimura varieties

### v1

@@ 3.2.2 The case of nodal curves
  1. 1.
    $Q$ has length one at each prenode. In particular, there is a natural inclusion $\nu^{*}N_{X/Y}\subset N_{X^{\prime}/Y}(D)$.
  2. 2.
    Via this inclusion, $M_{X^{\prime}/Y}=\nu^{*}N_{X/Y}$.
@@
We may prove the claim étale locally, and therefore assume we have coordinates $y_{1},\ldots,y_{n}$ of $Y$ and that $X$ has ideal $(y_{1}y_{2},y_{3},\ldots,y_{n})$. On the branch cut out by $(y_{2},y_{3},\ldots,y_{n})$, $N^{\vee}_{X^{\prime}/Y}$ is generated by $dy_{2},\ldots,dy_{n}$, while $\nu^{*}N^{\vee}_{X/Y}$ is generated by $y_{1}dy_{2},dy_{3},\ldots,dy_{n}$. This gives the first claim. Since $M_{X^{\prime}/Y}$ is generated by $y_{1}^{-1}\frac{\partial}{\partial y_{2}},\frac{\partial}{\partial y_{3}},\ldots,\frac{\partial}{\partial y_{n}}$, the second claim is established as well. ∎
Let $W_{X^{\prime}/Y}$ be the subsheaf of $(\iota^{\prime*}T_{Y})(D)$ of sections with simple poles at each prenode in the tangent direction of the attached branch. By the lemma, there is then a natural exact sequence
$0\to\nu^{*}N_{X/Y}^{\vee}\to W^{\vee}_{X^{\prime}/Y}\to\Omega_{X^{\prime}}\to 0$ (3.2.2)
Choose a splitting of (3.2.2) on each $X_{i}^{\prime}$ and use it to obtain $\hat{t}_{i}$ such that the following diagram commutes.
[…]
The lifts $\hat{t}_{i}$ yield $\eta_{i}:F^{*}\Omega_{X_{i}^{\prime}}\to\Omega_{X_{i}^{\prime}}$ such that $\eta_{i}\circ F^{*}\rho_{Y_{i}|X^{\prime}_{i}}=\rho_{Y_{i}|X^{\prime}_{i}}\circ\iota^{\prime*}\zeta_{i}-d\hat{t}_{i}$ as maps $F^{*}W_{X^{\prime}_{i}/Y_{i}}^{\vee}\to\Omega_{X_{i}^{\prime}}$, and moreover on overlaps we may write $u_{ij}\circ F^{*}\rho_{Y_{ij}|X^{\prime}_{ij}}=\iota^{\prime*}s_{ij}-(\hat{t}_{i|j}-\hat{t}_{j|i})$ for $u_{ij}:F^{*}\Omega_{X_{ij}}\to\mathcal{O}_{X_{ij}}$. We have $du_{ij}=\eta_{i|j}-\eta_{j|i}$.
Let $(V,\theta)\in HIG_{m}(Y|X^{\prime})$ be a pullback Higgs sheaf. Let $J$ be the ideal of the (reduced) set of prenodes of $X^{\prime}$ at which $\theta$ is nonzero in the tangent direction of the attached branch. We write $\hat{t}_{i}(F^{*}\theta)$ for the resulting composition
$F^{*}(J\otimes V|_{X_{i}^{\prime}})\xrightarrow{F^{*}\theta}F^{*}(V|_{X_{i}^{\prime}}\otimes W_{X_{i}^{\prime}/Y_{i}}^{\vee})\xrightarrow{\operatorname{id}\otimes\hat{t}_{i}}F^{*}V|_{X_{i}^{\prime}}$
and likewise for $\iota^{\prime*}s_{ij}(F^{*}\theta)$ and $u_{ij}\circ F^{*}\rho_{Y|X^{\prime}}(F^{*}\theta)$. We therefore have local morphisms
$\exp(-\hat{t}_{i}(F^{*}\theta)):F^{*}(J^{m-1}\otimes V)|_{X_{i}^{\prime}}\to V|_{X_{i}^{\prime}}$
and, because the overlaps contain no nodes, commutative diagrams on the overlaps
[…]
Denote by $(V,\theta_{X})$ the Higgs sheaf associated to $(V,\theta)$ and let $(G,0)\subset(V,\theta_{X})$ be a Higgs subsheaf with vanishing Higgs field. By restricting the above local endomorphisms, we obtain local maps
$\exp(-\hat{t}_{i}(F^{*}\theta)):F^{*}(J^{m-1}\otimes G)|_{X_{i}^{\prime}}\to F^{*}V|_{X_{i}^{\prime}}$ (3.2.3)
and commutative diagrams on the overlaps
[…]
again because there are no nodes in the overlaps.
@@
The local maps (3.2.3) give morphisms of flat sheaves
$\exp(-\hat{t}_{i}(F^{*}\theta)):(F^{*}(J^{m-1}\otimes G)|_{X_{i}^{\prime}},\nabla_{can})\to(F^{*}V|_{X_{i}^{\prime}},\nabla_{can}+\rho_{Y_{i}|X^{\prime}_{i}}\circ\iota^{\prime*}\zeta_{i}(F^{*}\theta))$
and glue via (3.2.4) to a global morphism of flat sheaves $(F^{*}(J^{m-1}\otimes G),\nabla_{can})\to C_{\tilde{Y}|X^{\prime}}^{-1}(V,\theta)$. The morphism does not depend on the choice of splittings of (3.2.2).
@@
We first show that the local map (3.2.3) respects the connections. Let $s$ be a local section of $J^{m-1}\otimes G$, so that $\nabla_{can}F^{*}s=0$. We have
$\nabla_{can}\exp(-\hat{t}_{i}(F^{*}\theta))F^{*}s$ $=\exp(-\hat{t}_{i}(F^{*}\theta))(-d\hat{t}_{i}(F^{*}\theta))F^{*}s$
$=\exp(-\hat{t}_{i}(F^{*}\theta))(-\rho_{Y_{i}|X^{\prime}_{i}}\circ\iota^{\prime*}\zeta_{i}(F^{*}\theta))F^{*}s$
$=-\rho_{Y_{i}|X_{i}^{\prime}}\circ\iota^{\prime*}\zeta_{i}(F^{*}\theta)\exp(-\hat{t}_{i}(F^{*}\theta))F^{*}s$
which implies the compatibility of the connections. According to the description preceding Lemma 3.2, these maps glue as morphisms of flat sheaves since the vertical map of (3.2.4) is the gluing map for $C_{\tilde{Y}|X^{\prime}}^{-1}(V,\theta)$.
A different choice of splittings of (3.2.2) yields $\hat{t}_{i}^{\prime}$ for which $\hat{t}_{i}-\hat{t}_{i}^{\prime}=r_{i}\circ F^{*}\rho_{Y_{i}|X^{\prime}_{i}}$ for $r_{i}:F^{*}\Omega_{X_{i}^{\prime}}\to\mathcal{O}_{X_{i}^{\prime}}$. Since $r_{i}\circ F^{*}\rho_{Y_{i}|X^{\prime}_{i}}(F^{*}\theta)$ kills $F^{*}(J^{m-1}\otimes G)$, the maps in (3.2.3) do not depend on the choice of lift. ∎
Remark 3.6.
Its reasonable to ask whether there is a version of Lemma 3.5 for a Higgs subsheaf $(G,\psi)\subset(V,\theta_{X})$ for which the Higgs field is nonzero. The maps (3.2.3) glue via the maps $\exp(u_{ij}(F^{*}\psi))$, and the natural local connection which glues is $\nabla_{can}+\eta_{i}(F^{*}\psi)$, which must be interpreted as a connection with log poles. It is no longer clear that either the resulting flat sheaf with log poles or the morphism is independent of the choice of splittings of (3.2.2) (even up to isomorphism).
In summary, we have the following:
Proposition 3.7.
Let $\tilde{Y}/W_{2}$ be smooth, $\tilde{X}/W_{2}$ a nodal curve, $\tilde{f}:\tilde{X}\to\tilde{Y}$ a $W_{2}$-morphism, $\nu:X^{\prime}\to X$ the normalization of the fiber over $k$, and $f^{\prime}=f\circ\nu$. Let $(V,\theta)\in HIG_{m}(Y)$ with $m\leq p$, and let $J$ be the ideal of the (reduced) set of prenodes in $X^{\prime}$ at which $\theta$ is nonzero in the tangent direction of the attached branch. Then for any Higgs subsheaf $(G,0)\subset f^{\prime*}(V,\theta)$ there is a flat subsheaf $(F^{*}(J^{m-1}\otimes G),\nabla_{can})\subset f^{\prime*}C^{-1}_{\tilde{Y}}(V,\theta)$. The morphism of the underlying sheaves is locally isomorphic to the natural inclusion $F^{*}(J^{m-1}\otimes G)\subset F^{*}f^{\prime*}V$.
@@
By [OV], the inverse Cartier transform commutes with pullback under morphisms of smooth $k$-varieties which lift to $W_{2}$. Thus we may assume $f$ is an embedding $\iota$ by replacing $Y$ with $\mathbb{P}^{N}\times Y$. The claim then follows from Lemma 3.2 and Lemma 3.5. ∎
3.3 Applications to Fontaine–Laffaille modules
In this subsection we prove some degree bounds on Higgs subsheaves of the graded object of a Fontaine–Laffaille module pulled-back to a nodal curve (Propositions 3.9 and 3.10 below). Recall [OV, §4.6] that we may think of a (strict) $p$-torsion Fontaine–Laffaille module $\mathbf{M}=(M,\nabla,F^{\bullet},\Phi)$ on a lift $\tilde{Y}$ of a smooth variety $Y/k$ as a locally free flat sheaf $(M,\nabla)$ with a locally split filtration $F^{\bullet}$ satisfying Griffiths transversality and equipped with an isomorphism of flat sheaves
$\Phi:C^{-1}_{\tilde{Y}}\operatorname{gr}_{F}\mathbf{M}\xrightarrow{\cong}(M,\nabla).$
where $\operatorname{gr}_{F}\mathbf{M}:=(\operatorname{gr}_{F}M,\operatorname{gr}_{F}\nabla)$ is the associated Higgs bundle. The category of $p$-torsion Fontaine–Laffaille modules $\mathbf{M}$ whose graded object $\operatorname{gr}_{F}M$ has nonzero pieces in degrees $[a,w+a]$ is denoted $\mathrm{MF}^{\nabla}_{[a,w+a],1}(\tilde{Y})$. Note that for $p$-torsion Fontaine–Laffaille modules, the index of the Hodge filtration may be freely shifted, as there is no integrality to force the degrees to be nonnegative.
We give the setup for these results now for convenience.
Setup 3.8.
Let $\tilde{Y}/W_{2}$ be a smooth $W_{2}$-scheme equipped with a strict $p$-torsion Fontaine–Laffaille module $\mathbf{M}=(M,\nabla,F^{\bullet}M,\Phi)\in\mathrm{MF}^{\nabla}_{[a,w+a],1}(\tilde{Y})$. Assume $p$ is an odd prime with $p-1\geq w$.
Let $\tilde{X}/W_{2}$ be a nodal curve together with a $W_{2}$-morphism $\tilde{f}:\tilde{X}\to\tilde{Y}$. Let $\nu_{B}:B\to X$ be the normalization of a proper component of the special fiber $X$ and let $g:B\to Y$ be the composition $g=f\circ\nu_{B}$. Let $D_{B,\mathbf{M}}$ the divisor of prenodes on $B$ at which the Higgs field $\operatorname{gr}_{F}\nabla$ is nonzero in the tangent direction of the attached branch, and $d_{B,\mathbf{M}}=\deg D_{B,\mathbf{M}}$ its degree.
Proposition 3.9.
Assume Setup 3.8. Suppose $(V,\theta)\subset g^{*}\operatorname{gr}_{F}\mathbf{M}$ is a Higgs subsheaf. Then $\mu(V)\leq wd_{B,\mathbf{M}}$ provided $p-1>wd_{B,\mathbf{M}}\operatorname{rk}(V)$.
Proof.
We argue as in the proof of Ogus–Vologodsky [OV]. Note that the exponent of $\operatorname{gr}_{F}\mathbf{M}$ is $m\leq w+1$. We first reduce to the case $(V,\theta)=(L,0)$ with rank one by taking the $\operatorname{rk}(V)$-th wedge power. We must then show that any Higgs sheaf $(L,0)\subset g^{*}\operatorname{gr}_{F}\mathbf{M}$ with $\operatorname{rk}(L)=1$ and vanishing Higgs field has $\deg(L)\leq wd_{B,\mathbf{M}}$. Suppose this is not the case, and let $(L,0)$ be a rank one Higgs subsheaf with vanishing Higgs field of maximal degree, $\deg(L)>wd_{B,\mathbf{M}}$. Let $J_{B,\mathbf{M}}=\mathcal{O}_{B}(-D_{B,\mathbf{M}})$. By Proposition 3.7 we have that
$(F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L),\nabla_{can})\to g^{*}C^{-1}_{\tilde{Y}}\operatorname{gr}_{F}\mathbf{M}\xrightarrow{g^{*}\Phi}g^{*}(M,\nabla)$
is an injective morphism of flat sheaves. Let $i\geq 0$ be minimal such that
$F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L)\to g^{*}M/F^{i+1}g^{*}M$
is nonzero. We therefore have an injective map
$F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L)\to\operatorname{gr}^{i}_{F}g^{*}M.$ (3.3.1)
Since $(F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L),\nabla_{can})$ is a flat subsheaf and the Higgs field of $\operatorname{gr}_{F}g^{*}M$ is $\operatorname{gr}_{F}\nabla$, (3.3.1) yields an injective morphism of Higgs sheaves
$(F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L),0)\to g^{*}\operatorname{gr}_{F}\mathbf{M}.$
But
$\deg(F^{*}(J_{B,\mathbf{M}}^{m-1}\otimes L))$ $=p(\deg(L)-(m-1)d_{B,\mathbf{M}})$
$\geq p(\deg(L)-wd_{B,\mathbf{M}})$
$\geq\deg(L)-wd_{B,\mathbf{M}}+(p-1)$
$>\deg(L)$
because of the assumption on $p$, and this contradiction finishes the proof. ∎
Let $(V,\theta)$ be a Higgs bundle on a smooth variety $Y/k$. There is a natural endomorphism Higgs bundle $(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)$. The Kodaira–Spencer map of $(V,\theta)$ is the Higgs field of $(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)$, thought of as a map $\operatorname{ad}\theta:T_{Y}\to\operatorname{\mathsf{End}}(V)$. For $\mathbf{M}\in\operatorname{MF}^{\nabla}_{[a,w+a],1}(Y)$, there is also a natural endomorphism object $\operatorname{\mathsf{End}}(\mathbf{M})=(\operatorname{\mathsf{End}}(M),\operatorname{ad}\nabla,F^{\bullet}\operatorname{\mathsf{End}}(M),\operatorname{ad}\Phi)\in\operatorname{MF}^{\nabla}_{[-w,w],1}(Y)$ in the category of Fontaine–Laffaille modules. The Kodaira–Spencer map of $\mathbf{M}$ is the Kodaira–Spencer map of the associated Higgs sheaf $\operatorname{gr}_{F}\mathbf{M}$.
Proposition 3.10.
Assume Setup 3.8 and let $P=\operatorname{img}(g^{*}\operatorname{gr}_{F}\operatorname{ad}\nabla)$ be the image of the pullback of the Kodaira–Spencer map $g^{*}\operatorname{gr}_{F}\operatorname{ad}\nabla:g^{*}T_{Y}\to g^{*}\operatorname{gr}_{F}\operatorname{\mathsf{End}}(\mathbf{M})$. Then $P$ does not contain a positive degree subsheaf of rank $r$ provided $p>2rwd_{B,\wedge^{r}\operatorname{\mathsf{End}}(\mathbf{M})}$.
Proof.
For simplicity of notation, set $(V,\theta)=\operatorname{gr}_{F}\mathbf{M}$. Observe that the Kodaira–Spencer map $\operatorname{ad}\theta:(T_{Y},0)\to(\operatorname{\mathsf{End}}(V),\operatorname{ad}\theta)$ is a morphism of Higgs sheaves, as therefore is $\wedge^{r}\operatorname{ad}\theta:(\wedge^{r}T_{Y},0)\to(\wedge^{r}\operatorname{\mathsf{End}}(V),\wedge^{r}\operatorname{ad}\theta)$. Thus, we have morphisms of flat sheaves
$(F^{*}\wedge^{r}T_{Y},\nabla_{can})=C_{\tilde{Y}}^{-1}(\wedge^{r}T_{Y},0)\to C_{\tilde{Y}}^{-1}(\wedge^{r}\operatorname{\mathsf{End}}(V),\wedge^{r}\operatorname{ad}\theta)\xrightarrow{\Phi}(\wedge^{r}\operatorname{\mathsf{End}}(M),\wedge^{r}\operatorname{ad}\nabla)$
and so morphisms of flat sheaves
$g^{*}(F^{*}\wedge^{r}T_{Y},\nabla_{can})\to g^{*}C_{\tilde{Y}}^{-1}\operatorname{gr}_{F}\wedge^{r}\operatorname{\mathsf{End}}(\mathbf{M})\xrightarrow{g^{*}\Phi}g^{*}(\wedge^{r}\operatorname{\mathsf{End}}(M),\wedge^{r}\operatorname{ad}\nabla).$
Observe that the image of the first map is naturally identified with $(F^{*}\wedge^{r}P,\nabla_{\mathrm{can}})$. Indeed, its enough to check this on the level of sheaves. The first map is locally identified with $g^{*}F^{*}\wedge^{r}\operatorname{ad}\theta$, while $(F^{*}\wedge^{r}P,\nabla_{\mathrm{can}})$ is the image of $F^{*}\wedge^{r}g^{*}\operatorname{ad}\theta$, since $F^{*},\wedge^{r}$ preserve injections and surjections of locally free sheaves. The functors $F^{*},g^{*},\wedge^{r}$ all naturally commute, so the claim follows.
If $G\subset P$ is subsheaf with $\operatorname{rk}(G)=r$, then setting $L=\wedge^{r}G$ by the same argument as in the previous proof we get an injective morphism of Higgs sheaves
$(F^{*}L,0)\to g^{*}\operatorname{gr}_{F}\wedge^{r}\operatorname{\mathsf{End}}(\mathbf{M})$
but then by Proposition 3.9 we would have
$2rwd_{B,\wedge^{r}\operatorname{\mathsf{End}}(\mathbf{M})}\geq\deg(F^{*}L)=p\deg(L)$
which given the assumption on $p$ means we must have $\deg(L)=\deg(G)\leq 0$.
Remark 3.11.
The same proof shows for instance that in Setup 3.8, $g^{*}\operatorname{gr}^{0}_{F}M$ does not contain a positive degree subsheaf of rank $r$ provided $p>rwd_{B,\wedge^{r}\mathbf{M}}$.
Proposition 3.10 will allow us to control the liftability of rational components of nodal curves with boundedly many nodes. This will be the key ingredient for the surface case of Proposition 4.4. The normal case requires a bit more care, so we only treat the smooth case now, and postpone the general setup until then.
Definition 3.12.
Given a Higgs bundle $(V,\theta)$ on $Y$, we say the Kodaira–Spencer map is quasi-finite if there is a partition of $Y$ into locally closed smooth subschemes $Y^{i}$ on which the Kodaira–Spencer map $T_{Y^{i}}\to\operatorname{\mathsf{End}}(V)|_{Y^{i}}$ is immersive, meaning it is injective with locally-free cokernel.
Of course, if the Kodaira–Spencer map of $(V,\theta)$ is already immersive, then $(V,\theta)$ has quasi-finite Kodaira–Spencer map, and this will be the relevant case for Shimura varieties.
Corollary 3.13.
Assume Setup 3.8, and further assume that the Kodaira–Spencer map of $\operatorname{gr}_{F}(M,\nabla)$ is quasifinite. Then if $B$ is rational, $g:B\to Y$ is constant provided $p>2wd_{B,\operatorname{\mathsf{End}}(\mathbf{M})}$
Proof.
[…]
Corollary 3.14.
Assume Setup 3.8, and further assume that the Kodaira–Spencer map of $\operatorname{gr}_{F}(M,\nabla)$ is quasifinite. Finally, assume either
  1. 1.
    $\tilde{X}$ is obtained by successively blowing up a smooth curve $\tilde{X}^{\prime}/W_{2}$ at closed points and $p>2w$.
  2. 2.
    $\tilde{X}$ is obtained by successively blowing up a nodal curve $\tilde{X}^{\prime}/W_{2}$ at closed points and $p>4w$.
Then $X\to Y$ factors through $X\to X^{\prime}$.
Proof.
The dual graph of the exceptional curves is a tree and the components that are not contracted form a subtree. The extremal components have $d_{B,\mathbf{M}}\leq 1$ in the first case and $\leq 2$ in the second, so the claim follows form the previous corollary. ∎

### v2

@@ 3.1.4 Local systems
Let $\mathcal{S}$ be a smooth formal $\mathcal{O}$-scheme, and $\mathcal{X}$ be either $\operatorname{Spf}\mathcal{O}^{\prime}$ or a smooth formal $\mathcal{O}$-scheme, and $f:X\rightarrow\mathcal{S}$. Let $\mathbf{M}\in\operatorname{MF}_{[0,w]}(\mathcal{S})$. Then we have a natural isomorphism $f^{-1}\mathbb{D}(M)\rightarrow\mathbb{D}(\operatorname{MF}(f)(M))$.
@@
Assume first that $\mathcal{X}$ is a smooth formal $\mathcal{O}$-scheme, and assume $\mathcal{X}=\operatorname{Spf}R,\mathcal{S}=\operatorname{Spf}A$ are both small and affine. Now we have a natural map $B^{+}(A)\rightarrow B^{+}(R)$ compatible with the Galois actions. Moreover, since $M$ is a vector bundle, we have that $\overline{C}_{\mathbf{M}}(B^{+}(A))\otimes B^{+}(R)\cong\overline{C}_{\mathbf{M}}(B^{+}(R))$. We therefore obtain a natural morphism $f^{-1}\mathbb{D}(\mathbf{M})\rightarrow\mathbb{D}(\operatorname{MF}(f)(\mathbf{M}))$. It remains to show that this is an isomorphism.
By [faltings, II,h] we see that the natural map $\overline{C}_{\mathbf{M}}(B^{+}(A))\rightarrow\mathbb{D}(\mathbf{M})^{*}\otimes B^{+}(A)$ is an isomorphism up to $\beta_{0}^{a}$ where $\beta_{0}$ is as in [faltings]. It follows that
$\beta_{0}^{a}\cdot\mathbb{D}(\operatorname{MF}(f)(\mathbf{M}))\otimes B^{+}(R)\subset f^{-1}\mathbb{D}(\mathbf{M})\otimes B^{+}(R)\subset\mathbb{D}(\operatorname{MF}(f)(\mathbf{M}))\otimes B^{+}(R).$
Finally, since $p\nmid\beta_{0}^{a}$, it follows that the map $f^{-1}\mathbb{D}(\mathbf{M})\rightarrow\mathbb{D}(\operatorname{MF}(f)(\mathbf{M}))$ is an isomorphism as desired. Since this map is clearly natural, we may glue up to obtain the claim for a general smooth formal $\mathcal{O}_{0}$ schemes $X,S$.
If $X=\operatorname{Spf}\mathcal{O}^{\prime}$ the proof is identical, using $B^{+}(\mathcal{O}^{\prime})$ instead of $B^{+}(R)$ and [faltingsramified, Thm 5] instead of [faltings, II,h],
3.1.5 Liu–Zhu correspondence
Let $\mathscr{X}$ denote a smooth geometrically connected rigid analytic variety over a finite extension $F^{\prime}/\mathbb{Q}_{p}$. Let $\mathbb{L}$ denote a de Rham $\mathbb{Q}_{p}$-local system on the etale site $\mathscr{X}_{{\mathrm{et}}}$. Then [liuzhu, Thm 3.9] associate a filtered flat bundle with Griffiths transverse connection $LZ(\mathbb{L})$ in a functorial way with respect to both $\mathscr{X}$ and $\mathbb{L}$. If $\mathscr{X}=\mathcal{X}^{rig}$ for a smooth, small formal scheme $\mathcal{X}=\operatorname{Spf}R$, then $LZ(\mathbb{L})\cong(\mathbb{L}\otimes\mathcal{O}B_{{}_{\mathrm{dR}}}(R[1/p]))^{G_{R}}$.
@@ 3.1.5 Liu–Zhu correspondence
Let $\mathcal{X}$ denote either a smooth formal scheme over $\mathcal{O}$ or $\operatorname{Spf}\mathcal{O}^{\prime}$. Let $\mathbf{M}\in\operatorname{MF}_{[0,w]}(\mathcal{X})$. Then there is a canonical isomorphism
$\begin{cases}M[1/p]\rightarrow LZ(\mathbb{D}(M)^{*})&\textrm{ if }\mathcal{X}\textrm{ is smooth over }\mathcal{O}\\ M\otimes_{R_{\mathcal{O}^{\prime}}}F^{\prime}\rightarrow LZ(\mathbb{D}(M)^{*})&\textrm{ if }\mathcal{X}=\operatorname{Spf}\mathcal{O}^{\prime}\\ \end{cases}$
@@
We once again first work locally on $\operatorname{Spf}R$ for a small $R$. By [liuzhu, Thm 3.9,v], we see that
$LZ(\mathbb{D}(\mathbf{M})^{*})\otimes_{R}\mathcal{O}B_{{}_{\mathrm{dR}}}(R[1/p])\rightarrow\mathbb{D}(\mathbf{M})^{*}\otimes\mathcal{O}B_{{}_{\mathrm{dR}}}(R[1/p]))$
is an isomorphism
Moreover, there is a natural map [Brinon] $B^{+}(R)[\beta_{0}^{-1}]\rightarrow\mathcal{O}B_{{}_{\mathrm{dR}}}(R[1/p])$. Hence, composing the above with [faltings, II, Theorem 2.6*, h] we see that
$LZ(\mathbb{D}(\mathbf{M})^{*})\otimes\mathcal{O}B_{{}_{\mathrm{dR}}}(R[1/p])\cong C_{\mathbf{M}}(B^{+}(R))\otimes_{B^{+}(R)}\mathcal{O}B_{{}_{\mathrm{dR}}}(R[1/p]).$
Now $\mathcal{O}B_{{}_{\mathrm{dR}}}(R[1/p])$ maps to $\widehat{\overline{R}}[1/p]$, we obtain a Galois-equivariant morphism
$LZ(\mathbb{D}(\mathbf{M})^{*})\otimes\hat{\overline{R}}[1/p]\cong\overline{C}_{\mathbf{M}}(B^{+}(R))\otimes_{B^{+}(R)}\hat{\overline{R}}[1/p]\cong\overline{C}_{\mathbf{M}}(\widehat{\overline{R}})[1/p]\cong M\otimes_{R}\widehat{\overline{R}}[1/p].$
Taking Galois invariants, we see that $LZ(\mathbb{D}(\mathbf{M})^{*})\cong M[1/p]$. Since all the maps are natural, this glues up to completes the proof for an arbitrary smooth formal scheme $\mathcal{O}$.
$LZ(\mathbb{D}(\mathbf{M})^{*})\cong\left(\overline{C}_{\mathbf{M}}(B^{+}(R))\otimes_{B^{+}(R)}\mathcal{O}B_{{}_{\mathrm{dR}}}(R[1/p])\right)^{G_{R}}\rightarrow\overline{C}_{\mathbf{M}}(\widehat{\overline{R}})[1/p]^{\operatorname{Gal}(R)}\cong(M\otimes_{R}\widehat{\overline{R}}[1/p])^{\operatorname{Gal}(R)}\cong M[1/p]$
For $X=\mathcal{O}^{\prime}$ the proof is exactly the same, noting that in this case for $\mathbf{M}\in\operatorname{MF}_{[0,w]}(\mathcal{O}^{\prime})$, we have $C_{\mathbf{M}}(R)\cong M\otimes_{R_{\mathcal{O}^{\prime}}}\mathcal{O}^{\prime}$. ∎
3.2 Proof of Theorem 3.1
We use the notion of crystallinity in [guo, Def 4.4], which we refer to as Brinon-crystalline. Recall that a classical point of a rigid space $\mathscr{Y}$ over $F$ is a map $\operatorname{spa}F^{\prime}\rightarrow\mathscr{Y}$ for a finite extension $F^{\prime}/F$.
Lemma 3.8.
Let $\mathcal{O}$ be the ring of integers in a finite unramified extension $F/\mathbb{Q}_{p}$, and let $\mathcal{X}/\operatorname{Spf}\mathcal{O}$ be a smooth formal scheme. Let $\mathbb{L}$ be a $\mathbb{Z}_{p}$-local system on $\mathcal{X}^{rig}$ which is Brinon-crystalline at every classical point of $\mathcal{X}^{rig}$. Then $\mathbb{L}$ is crystalline.
@@
By [guo, Thm 1.1] we obtain that $\mathbb{L}$ is Brinon-crystalline. Since $F/\mathbb{Q}_{p}$ is unramified, it follows by [Hokaj, Thm 3] that $\mathbb{L}$ is crystalline. ∎
Proof of Theorem 3.1.
Let $\mathbb{M}$ be the $\mathbb{Z}_{p}$-local system on $\mathcal{S}^{rig}$ corresponding to $\mathbf{M}$ on $\hat{\mathcal{S}}$, and $\mathbb{N}=(f^{rig})^{*}\mathbb{M}$ the pullback. By Lemma 3.8, $\mathbb{N}$ is crystalline, hence corresponds to a Fontaine–Laffaille module $\mathbf{N}$ on $\hat{\mathcal{X}}$. Now by Lemma 3.7 we see that we have maps $r_{1}:g^{*}M\rightarrow LZ(g^{-1}\mathbb{D}(\mathbf{M})^{*})$ and $r_{2}:\pi^{*}N\rightarrow LZ(\pi^{-1}\mathbb{D}(\mathbf{N})^{*})$, and natural isomorphisms $g^{-1}D(\mathbf{M})^{*}\rightarrow\pi^{-1}D(\mathbf{N})^{*}$ giving induced isomorphisms
$s:LZ(g^{-1}\mathbb{D}(\mathbf{M})^{*})\rightarrow LZ(\pi^{-1}\mathbb{D}(\mathbf{N})^{*}).$
We claim that the images of the maps $r_{1},r_{2}$ are identified by $s$. To prove this its enough to work at every $\overline{\mathbb{Q}_{p}}$-point. So let $F^{\prime}/\mathbb{Q}_{p}$ be a local field and $\mathcal{O}^{\prime}$ its ring of integers and $\iota:\operatorname{Spf}\mathcal{O}^{\prime}\rightarrow\mathcal{X}^{\prime}$ an arbitrary point. Note that by Lemma 3.6 we have that $\operatorname{MF}(g\circ\iota)(\mathbf{M})$ is a Fontaine-Laffaille module, and $\mathbb{D}(\operatorname{MF}(g\circ\iota)(\mathbf{M}))\cong\iota^{-1}g^{-1}\mathbb{D}(\mathbf{M})$, and that $\mathbb{D}(\operatorname{MF}(g\circ\iota)(\mathbf{N}))\cong\iota^{-1}\pi^{-1}\mathbb{D}(\mathbf{N})$.
By Lemmas 3.6, 3.7 we obtain maps
$\iota^{*}r_{1}:\iota^{*}g^{*}M\rightarrow\iota^{*}LZ(g^{-1}\mathbb{D}(\mathbf{M})^{*})\cong LZ(\mathbb{D}(\operatorname{MF}(g\circ\iota)(\mathbf{M}))^{*}).$
Moreover, this map is canonical and therefore so is its image. It follows that the images of $\iota^{*}r_{1}$ and $\iota^{*}r_{2}$ agree. Thus, the images of $r_{1}$ and $r_{2}$ agree, as desired. It follows that $g^{*}M$ and $\pi^{*}M$ are isomorphic as filtered flat bundles, and thus the Griffiths bundle of $g^{*}(M,\nabla,F^{\bullet}M)$ is trivial on $\pi^{-1}(x)$, as desired.
It follows by Lemma 2.17 that if the Griffiths bundle of $(M,\nabla,F^{\bullet}M)$ is ample, the map $g$ descends to $\mathcal{X}$, completing the proof.